Stock Y has a beta of 1.30 and an expected return of 13.5 percent. Stock Z has a beta of .75 and an expected return of 10.6 percent. If the risk-free rate is 4.75 percent and the market risk premium is 7.25 percent, are these stocks overvalued or undervalued?

stock Y = ______

stock Z = ______
Answer #1

Expected return = Risk-free rate + Beta(Market risk premium)

Stock Y = 0.0475 + 1.30(0.0725) = 0.1418 or 14.18%

Stock Y is overvalued.

Stock Z = 0.0475 + 0.75(0.0725) = 0.1019 or 10.19%

Stock Z is undervalued.